# Multivalent linkers mediated ultra-sensitive bio-detection

### v2

@@ Abstract
In biosensing and diagnostic applications, a key objective is to design detection systems capable of identifying targets at very low concentrations, i.e., achieving high sensitivity. Here, we propose a linker-mediated detection scheme in which the presence of multivalent target molecules (linkers) facilitates the adsorption of ligand-coated guest nanoparticles onto a receptor-coated host substrate. Through a combination of computer simulations and mean-field theory, we demonstrate that, at fixed overall binding strength, increasing the valency of linkers exponentially lowers the concentration threshold for detection. This counterintuitive behavior arises from the combinatorial entropy associated with multivalent binding configurations, which tremendously amplifies the adsorption sensitivity and enables the identification of targets at extremely low concentrations. Our findings highlight multivalency engineering of linkers as a powerful strategy to substantially enhance the sensitivity of biodetection systems.
@@
Introduction
Biomarker and pathogen detections are pivotal for early-stage biomarker and pathogen diagnostics Shaw et al. (2007); Jain and Jain (2021). Yet conventional platforms such as ELISA and PCR often require labor-intensive sample preparation and specialized infrastructure, which hampers rapid, low-cost deployment in complex sample matrices Engvall and Perlmann (1971); Mullis and Faloona (1987); Aydin (2015); Bustin and Nolan (2004). A central challenge is to obtain ultrasensitive, superselective, and fast readouts at physiologically low analyte concentrations Wulfkuhle et al. (2003). Multivalency offers an important route to this goal: by forming multiple, individually weak bonds, a probe achieves strong adsorption and sharp, nonlinear responses to target abundance Kitov and Bundle (2003); Huskens (2006); Martinez-Veracoechea and Frenkel (2011). Beyond the classical picture of multivalent guests adsorbing on receptor-coated host substrates Martinez-Veracoechea and Frenkel (2011), many experimentally relevant assays use linker-mediated architectures, in which soluble target molecules (linkers) bridge a ligand-coated guest to a receptor-decorated host. Prominent examples include genome detection using DNA-functionalized colloids whose sticky ends and mobile linkers enable “sandwich” binding, antibodies that bridge antigens and Fc-family receptors, e.g., via engineered bispecific or multivalent formats, and glyco-functionalized nanoparticles engaging lectin receptors on cell membranes Nimmerjahn and Ravetch (2008); Labrijn et al. (2019); Brinkmann and Kontermann (2017); Kiessling and Grim (2013); Wrapp et al. (2020); Budhadev et al. (2020); Curk et al. (2022); Xu et al. (2023); Xia and Ni (2024). However, the design principle of multivalent linkers remains unknown for biosensing. In particular, how does distributing a fixed total binding strength over more linker arms reshape the detection threshold? Here, building on a Langmuir-like adsorption framework Martinez-Veracoechea and Frenkel (2011) and the recent statistical-mechanical theory for multivalent bridging Xia et al. (2025), we develop a linker-mediated detection scheme that achieves both ultrasensitivity and superselectivity.
Model
@@ Introduction
To validate our theoretical prediction and provide a parameter mapping for future studies, we performed Grand Canonical Monte Carlo (GCMC) simulations of guest nanoparticles in contact with the host substrate coated with explicit receptors. Due to the low density and the multivalent nature of guest particles in most experiments, the equilibration time for 3D simulations is extremely long. Therefore, we simulate the a monolayer of guest particles on the substrate where the guest particles can be modelled as 2D hard disks, and they are in chemical equilibrium with a bulk 3D reservoir above. The adsorption and desorption of guest particles are treated as the insertion and deletion of particles from the reservoir in the simulation Supplementary Materials . Guest particles are modeled as hard disks with diameter $\sigma=1$ on a two-dimensional substrate in chemical equilibrium with a reservoir at chemical potential $\mu_{g}$ Xia et al. (2023); Xia and Ni (2024) (see Fig. 1b). The activity of guest particles is given by $z_{g}=e^{\beta\mu_{g}}\pi\sigma^{2}h_{0}/(4\Lambda^{3})$, where $h_{0}=1$ is the thickness of the adsorption layer. Immobile receptors are randomly distributed on the substrate following a Poisson distribution with number density $\rho_{r}$. Mobile ligands on each guest particle can interact with receptors located within the area underneath the particle. Guest particles move on a two-dimensional plane of box length $L$ with periodic boundary conditions and interact via hard-disk repulsion. The maximum number of adsorption sites is set to $N_{\rm max}=4L^{2}/(\pi\sigma^{2})$, and the adsorption probability is calculated as $\theta=\langle N_{g}\rangle/N_{\rm max}$, where $\langle N_{g}\rangle$ is the average number of adsorbed guest particles.
@@
Ultrasensitive $\rho$-response
To investigate how linker valency $\kappa$ affects the adsorption of guest nanoparticles, we perform both theoretical analysis and GCMC simulations, considering symmetric linkers with $\kappa=\kappa_{l}=\kappa_{r}$. We fix the overall binding strength to isolate the effect of linker valency so that any change in the adsorption of nanparticles arises solely from the combinatorial entropy associated with multivalency. Accordingly, in all calculations we fix the total binding strength as $f_{\rm tot}=\kappa f_{l}=\kappa f_{r}$.
@@
In addition to selectivity, linker valency also strongly influences the adsorption and desorption sensitivity thresholds, $\rho_{\rm ad}$ and $\rho_{\rm dp}$. As shown in Fig. 2d, $\rho_{\rm ad}$ decreases exponentially with increasing $\kappa$, whereas $\rho_{\rm dp}$ decreases and then reach a plateau at large $\kappa$. Since the overall binding strength is fixed, these trends arise from the growing combinatorial entropy introduced by linker multivalency. To elucidate the underlying physics, we examine the large-$\kappa$ limit at a single site.
@@
This shows that $\rho_{\rm ad}$ decreases exponentially with increasing linker valency $\kappa$, with the exponent given by the constant $\left[\log(\bar{n}_{l}+1)+\log(\bar{n}_{r}+1)\right]$ at the large $\kappa$ limit, which agrees quantitatively with GCMC simulations (Fig. 2d). Thus, as the total binding strength $\kappa f_{l/r}$ is fixed, this exponential scaling reflects the dominant contribution of the combinatorial entropy associated with linker multivalency. This intriguing result suggests that even with a fixed overall binding strength, increasing $\kappa$ can exponentially lower the adsorption threshold or detection limit, thereby enabling ultrasensitive detection.
Different from the effect of valency $\kappa$ on the adsorption of nanoparticles, the desorption threshold $\rho_{\rm de}$ decreases with increasing $\kappa$ and reaches a plateau at large $\kappa$ limit (Fig. 2d). This is because all receptors and ligands are fully occupied at large $\kappa$. Further increasing $\kappa$ does not increase the number of bound ends per linker, thus has no effect on the adsorption probability $\theta$ and corresponding $\rho_{\rm de}$ Supplementary Materials . This behavior is qualitatively consistent with the purely entropy-driven phase behavior observed in programmable colloidal atom-electron equivalents at large linker valency Xia et al. (2025).
Detection among non-specific binders
@@
In summary, we have developed a theoretical framework that elucidates how linker multivalency governs the sensitivity and selectivity of bio-detection systems. Our analysis shows that, at a fixed overall binding strength, increasing linker valency exponentially lowers the detection threshold, thereby enabling ultrasensitive readouts even at extremely low target concentrations. As a result, even sparse populations of multivalent targets can generate robust adsorption signals, which is highly desirable for early-stage biomarker and pathogen diagnostics. At the same time, increasing linker valency reduces superselectivity with respect to linker concentration, reflecting a fundamental trade-off between sensitivity and concentration-based discrimination. Furthermore, the presence of non-specific binders can suppress ultrasensitivity when their concentration becomes sufficiently high. This effect can be mitigated by increasing the densities of receptors grafted on host substrate or the number of ligands coated on guest nanoparticles, thereby restoring the detection sensitivity. Taken together, these findings indicate that optimal detection performance does not necessarily require ligands and receptors with the strongest possible binding affinities. Instead, those capable of forming multiple simultaneous interactions with the target may be more effective. This principle also rationalizes the recently developed ultrasensitive whole-genome detection strategy based on DNA-functionalized colloids, where long single-stranded DNAs (ssDNAs) act as “multivalent linkers”, driving colloidal aggregation at extremely low ssDNA concentrations Xu et al. (2023). Our results provide general design principles for engineering multivalent detection systems and highlight the fundamental role of combinatorial entropy in biomolecular recognition. The theoretical insights presented here offer practical guidelines for the development of next-generation biosensors and diagnostic assays capable of ultra-sensitive and highly specific detection in complex biological environments.
Acknowledgements.
@@
I Data Availability Statements

### v4

# Entropy-driven physical amplification in multivalent biosensing

@@ Abstract
Sensitive detection of low-abundance molecular targets is widely assumed to require enzymatic amplification, such as PCR, to achieve low detection limits. In amplification-free platforms, sensitivity is traditionally constrained by equilibrium binding affinity. Here we show that multivalent linker entropy provides a distinct physical route to exponential sensitivity enhancement in purely equilibrium sensing architectures. Using a statistical-mechanical theory supported by grand canonical Monte Carlo simulations, we demonstrate that redistributing a fixed total interaction strength over increasing linker valency exponentially lowers adsorption thresholds. This scaling emerges not from stronger energetic affinity, but from the rapid growth of combinatorial binding configurations, revealing entropy as an intrinsic amplification mechanism. Consequently, detection limits can be tuned independently of bond strength, enabling ultrasensitive responses without enzymatic replication. Our results establish a general physical design principle for engineering amplification-free detection systems capable of approaching PCR-level sensitivities through entropy-driven collective effects.
@@
I Introduction
Sensitive detection of biomarkers and pathogens requires achieving low detection limits at physiologically relevant analyte concentrations. In amplification-based platforms such as PCR, extraordinary sensitivity is obtained through exponential molecular replication Yang and Rothman (2004); Phaneuf et al. (2013). In contrast, amplification-free biosensing systems must rely directly on binding thermodynamics and collective molecular interactions to generate detectable signals de Picciotto et al. (2014); Vallee-Belisle et al. (2009). Understanding how such systems can approach amplification-level sensitivity without enzymatic replication remains a fundamental physical challenge Vallee-Belisle et al. (2009).
Multivalent interactions provide a general physical mechanism for generating nonlinear and highly cooperative responses in systems composed of many weak binding units Kitov and Bundle (2003); Huskens (2006a); Martinez-Veracoechea and Frenkel (2011). By distributing interactions over multiple binding sites, multivalency allows individually weak bonds to collectively stabilize macroscopic states, leading to sharp transitions in adsorption, adhesion, and molecular recognition Huskens (2006b). In the context of surface adsorption, statistical-mechanical frameworks have shown that multivalent binding can produce superlinear, and even superselective, responses to changes in environmental parameters Martinez-Veracoechea and Frenkel (2011). Importantly, existing studies have primarily focused on how multivalency sharpens response slopes, enabling discrimination based on receptor density or target abundance, whereas much less is known about how multivalency controls the location of response thresholds themselves, particularly under constraints on total interaction strength.
In many experimentally relevant systems, multivalent recognition does not occur through direct adsorption alone, but instead proceeds via linker-mediated architectures, in which soluble molecules bridge ligands on one object to receptors on another. Such bridging motifs are ubiquitous across bioanalytical and soft-matter platforms, including antibody-based sandwich immunoassays Engvall and Perlmann (1971); Voller et al. (1976), DNA-mediated nanoparticle assays Mirkin et al. (1996); Taton et al. (2000); Cao et al. (2002) and whole-genome detection Xu et al. (2023); Xia and Ni (2024). From a physical perspective, linker-mediated interactions fundamentally alter the structure of multivalent binding by introducing an additional degree of freedom: linkers can form productive bridges connecting the two surfaces, or remain non-productively sequestered by dangling from either side. This competition between bridging and sequestration gives rise to response behaviors that differ qualitatively from direct multivalent adsorption, including non-monotonic responses characterized by a low-concentration “turn-on” threshold and a high-concentration “turn-off” threshold that together define a finite operating window Tate and Ward (2004); Voller et al. (1976).
These features raise a fundamental question concerning multivalent interacting systems under constrained interaction strength. Specifically, when the total binding strength between two objects is fixed, what controls the position of the adsorption threshold? Is the threshold set primarily by energetic affinity, or can collective and entropic effects associated with multivalency qualitatively reshape it? In linker-mediated architectures, this question becomes particularly acute because linker valency provides a natural means of redistributing a fixed interaction strength over multiple binding arms, necessarily weakening individual bonds while expanding the space of accessible binding configurations, thereby introducing an intrinsic trade-off between energetic stabilization and combinatorial entropy.
Here we address this question by developing a minimal statistical-mechanical framework for linker-mediated multivalent adsorption. Combining a Langmuir-like adsorption description Martinez-Veracoechea and Frenkel (2011); Xia and Ni (2024), a theory for multivalent bridging Xia et al. (2025), and grand canonical Monte Carlo simulations, we systematically isolate the role of linker valency under the constraint of fixed total binding strength. We find that redistributing a fixed interaction strength over an increasing number of linker arms can exponentially shift the adsorption threshold to lower linker concentrations. This strong and counterintuitive scaling does not originate from enhanced energetic affinity, but instead from the rapid growth of the combinatorial entropy associated with multivalent binding configurations, revealing an entropy-dominated mechanism for threshold positioning.
II Results
II.1 Model and simulations
@@
To validate our theoretical prediction and provide a parameter mapping for future studies, we performed Grand Canonical Monte Carlo (GCMC) simulations of guest nanoparticles in contact with the host substrate coated with explicit receptors. Due to the low density and the multivalent nature of guest particles in most experiments, the equilibration time for 3D simulations is extremely long. Therefore, we simulate a monolayer of guest particles on the substrate where the guest particles can be modelled as 2D hard disks, and they are in chemical equilibrium with a bulk 3D reservoir above. The adsorption and desorption of guest particles are treated as the insertion and deletion of particles from the reservoir in the simulation Supplementary Materials . Guest particles are modeled as hard disks with diameter $\sigma=1$ on a two-dimensional substrate in chemical equilibrium with a reservoir at chemical potential $\mu_{g}$ Xia et al. (2023); Xia and Ni (2024) (see Fig. 1b). The activity of guest particles is given by $z_{g}=e^{\beta\mu_{g}}\pi\sigma^{2}h_{0}/(4\Lambda^{3})$, where $h_{0}=1$ is the thickness of the adsorption layer. Immobile receptors are randomly distributed on the substrate following a Poisson distribution with number density $\rho_{r}$. Mobile ligands on each guest particle can interact with receptors located within the area underneath the particle. Guest particles move on a two-dimensional plane of box length $L$ with periodic boundary conditions and interact via hard-disk repulsion. The maximum number of adsorption sites is set to $N_{\rm max}=4L^{2}/(\pi\sigma^{2})$, and the adsorption probability is calculated as $\theta=\langle N_{g}\rangle/N_{\rm max}$, where $\langle N_{g}\rangle$ is the average number of adsorbed guest particles.
@@
II.2 Ultrasensitive $\rho$-response
To investigate how linker valency $\kappa$ affects the adsorption of guest nanoparticles, we perform both theoretical analysis and GCMC simulations, considering symmetric linkers with $\kappa=\kappa_{l}=\kappa_{r}$. We fix the overall binding strength to isolate the effect of linker valency so that any change in the adsorption of nanoparticles arises solely from the combinatorial entropy associated with multivalency. Accordingly, in all calculations we fix the total binding strength as $f_{\rm tot}=\kappa f_{l}=\kappa f_{r}$.
@@ II.2 Ultrasensitive $\rho$-response
In addition to selectivity, linker valency also strongly influences the adsorption and desorption sensitivity thresholds, $\rho_{\rm ad}$ and $\rho_{\rm dp}$. As shown in Fig. 2d, $\rho_{\rm ad}$ decreases exponentially with increasing $\kappa$, whereas $\rho_{\rm dp}$ decreases and then reaches a plateau at large $\kappa$. Since the overall binding strength is fixed, these trends arise from the growing combinatorial entropy introduced by linker multivalency. To elucidate the underlying physics, we examine the large-$\kappa$ limit at a single site.
@@
This shows that $\rho_{\rm ad}$ decreases exponentially with increasing linker valency $\kappa$, with the exponent given by the constant $\left[\log(\bar{n}_{l}+1)+\log(\bar{n}_{r}+1)\right]$ at the large $\kappa$ limit, which agrees quantitatively with GCMC simulations (Fig. 2d). Thus, as the total binding strength $\kappa f_{l/r}$ is fixed, this exponential scaling reflects the dominant contribution of the combinatorial entropy associated with linker multivalency. This intriguing result suggests that even with a fixed overall binding strength, increasing $\kappa$ can exponentially lower the detection limit, thereby enabling ultrasensitive detection.
Different from the effect of valency $\kappa$ on the adsorption of nanoparticles, the desorption threshold $\rho_{\rm dp}$ decreases with increasing $\kappa$ and reaches a plateau at large $\kappa$ limit (Fig. 2d). This is because all receptors and ligands are fully occupied at large $\kappa$. Further increasing $\kappa$ does not increase the number of bound ends per linker, thus has no effect on the adsorption probability $\theta$ and corresponding $\rho_{\rm dp}$ Supplementary Materials . This behavior is qualitatively consistent with the purely entropy-driven phase behavior observed in programmable colloidal atom-electron equivalents at large linker valency Xia et al. (2025).
II.3 Detection among non-specific binders
@@ III Discussion
In summary, we have developed a theoretical framework that elucidates how linker multivalency governs the sensitivity and selectivity of bio-detection systems. Our analysis shows that, at a fixed overall binding strength, increasing linker valency exponentially lowers the detection threshold, thereby enabling ultrasensitive readouts even at extremely low target concentrations. In practical terms, amplification-free multivalent systems engineered in this manner may substantially narrow the sensitivity gap relative to PCR-based amplification assays. That such performance can arise purely from entropy-driven combinatorial effects, without enzymatic replication, highlights multivalency as a physical analogue of signal amplification. The exponential scaling identified here provides a mechanistic explanation for how ART-type platforms can achieve surprisingly high sensitivity despite lacking molecular amplification. As a result, even sparse populations of multivalent targets can generate robust adsorption signals, which is highly desirable for early-stage biomarker and pathogen diagnostics. At the same time, increasing linker valency reduces superselectivity with respect to linker concentration, reflecting a fundamental trade-off between sensitivity and concentration-based discrimination. Furthermore, the presence of non-specific binders can suppress ultrasensitivity when their concentration becomes sufficiently high. This effect can be mitigated by increasing the densities of receptors grafted on host substrate or the number of ligands coated on guest nanoparticles, thereby restoring the detection sensitivity. Taken together, these findings indicate that optimal detection performance does not necessarily require ligands and receptors with the strongest possible binding affinities. Instead, those capable of forming multiple simultaneous interactions with the target may be more effective. These principles are directly applicable to rapid antigen tests, DNA-mediated aggregation assays, and multivalent immunosensors, where signal generation relies on collective binding rather than enzymatic replication. This principle also rationalizes the recently developed ultrasensitive whole-genome detection strategy based on DNA-functionalized colloids, where long single-stranded DNAs (ssDNAs) act as “multivalent linkers”, driving colloidal aggregation at extremely low ssDNA concentrations Xu et al. (2023). Our results provide general design principles for engineering multivalent detection systems and highlight the fundamental role of combinatorial entropy in biomolecular recognition. The theoretical insights presented here offer practical guidelines for the development of next-generation biosensors and diagnostic assays capable of ultra-sensitive and highly specific detection in complex biological environments.
IV Method
In this work, grand-canonical Monte Carlo (GCMC) simulations are performed for 2D systems with implicit ligands and linkers. The guest particles are modelled as hard disks with radius $\sigma/2$, each coated with ligands. The ligands can interact with the receptors within $\sigma/2$ to the center of the guest particles. The receptors are distributed on the host substrate according to the Poisson distribution.
Due to the low density and the multivalent nature of guest particles in most experiments, the equilibration time for 3D simulations is extremely long. Therefore, we simulate the a monolayer of guest particles on the substrate where the guest particles can be modelled as 2D hard disks, and they are in chemical equilibrium with a bulk 3D reservoir above. The adsorption and desorption of guest particles are treated as the insertion and deletion of particles from the reservoir in the simulation. The probabilities of attempting translational trial moves, particle addition trial moves, and deletion trial moves are set to $0.3$, $0.35$, and $0.35$, respectively.
Moreover, if a particle is inserted at a random position $\bf s^{\prime}$ and its effective interaction is described by the adsorption partition function $q({\bf s^{\prime}})=e^{-\beta[F_{\rm ad}({\bf s^{\prime}})-F_{\rm de}({\bf s^{\prime}})]}-1$. This trial move is accepted by the probability
$acc\left(N_{g}\to N_{g}+1\right)=\\ {\rm min}\left[1,\frac{Vq({\bf s^{\prime}})}{\Lambda^{2}(N_{g}+1)}e^{\beta[\mu_{g}-U({\bf s^{\prime}},N_{g}+1)+U({\bf s^{\prime}},N_{g})]}\right],$ (8)
where $V$ is the area of the 2D plane, $\Lambda$ is the thermal de Broglie wavelength, $N_{g}$ is the number of guest particles before the insertion, $\beta\mu_{g}$ is the chemical potential of the reservoir of guest particles, and $U({\bf s^{\prime}},N_{g})$ quantifies the hard-core repulsion. The deletion of a particle at position $\bf s^{\prime}$ is accepted with probability
$acc\left(N_{g}\to N_{g}-1\right)=\\ {\rm min}\left[1,\frac{\Lambda^{2}N_{g}}{Vq({\bf s^{\prime}})}e^{-\beta[\mu_{g}+U({\bf s^{\prime}},N_{g}-1)-U({\bf s^{\prime}},N_{g})]}\right],$ (9)
where $N_{g}$ is the number of adsorbed guest particles before deletion. The translational trial move of a particle from a position $\bf s$ to $\bf s^{\prime}$ is accepted with a probability
$acc\left(\bf s\to s^{\prime}\right)={\rm min}\left[1,\frac{q({\bf s^{\prime}})}{q({\bf s})}e^{-\left(\beta U({\bf s^{\prime}},N_{g})-\beta U({\bf s},N_{g})\right)}\right].$ (10)
Moreover, periodic boundary conditions are applied in all directions. In each simulation, $10^{7}$ MC moves are performed for equilibration and $10^{8}$ MC moves for production with sampling every $5$ moves. The adsorption probability $\theta$ is computed as $N_{g}/N_{\rm max}$.
V acknowledgments
@@ IV Method
VI Data Availability Statements